# Efficacy of Language Model Self-Play in Non-Zero-Sum Games

### v1

@@ Abstract
Game-playing agents like AlphaGo have achieved superhuman performance through self-play, which is theoretically guaranteed to yield optimal policies in competitive games. However, most language tasks are partially or fully cooperative, so it is an open question whether techniques like self-play can effectively be used to improve language models. We empirically investigate this question in a negotiation game setting known as Deal or No Deal (DoND). Crucially, the objective in DoND can be modified to produce a fully cooperative game, a strictly competitive one, or anything in between. We finetune language models in self-play over multiple rounds of filtered behavior cloning in DoND for each of these objectives. Contrary to expectations, we find that language model self-play leads to significant performance gains in both cooperation and competition with humans, suggesting that self-play and related techniques have promise despite a lack of theoretical guarantees.
Efficacy of Language Model Self-Play in Non-Zero-Sum Games
Austen Liao   Nicholas Tomlin   Dan Klein Computer Science Division, University of California, Berkeley {austenliao, nicholas_tomlin, klein}@berkeley.edu
$**{}^{\text{*}}$qual contribution.
@@ 1 Introduction
Many of the greatest achievements in artificial intelligence have occurred in two-player zero-sum (2p0s) games such as Go (Silver et al., 2016), chess (Silver et al., 2018), and heads-up poker (Brown and Sandholm, 2018). One key technique enabling these breakthroughs has been self-play, in which identical copies of a model are pitted against each other and used to generate new training data. By iteratively training on their own data from games of self-play, models like AlphaGo and AlphaZero were able to continue improving long past the threshold of human performance. In certain types of 2p0s games, self-play is theoretically guaranteed to produce optimal policies, given sufficient model capacity and compute (Bai and Jin, 2020; Bai et al., 2020). However, in settings that involve collaboration with humans, self-play is no longer guaranteed to yield optimal policies (Strouse et al., 2021).
It is an open question whether the same principles that led to the success of models like AlphaGo can be applied to language models. Empirically, previous work on training agents to communicate via self-play has shown that they often invent uninterpretable communication strategies (Kottur et al., 2017); even when initialized with natural language data, self-play can cause models to gradually diverge from human-interpretable language (Lewis et al., 2017). As a result, much work has focused on mitigating these challenges, e.g., by regularizing with models trained on human data (FAIR, 2022).
@@
We then investigate the reasons behind these improvements, finding that models trained with self-play better follow task instructions, hallucinate less, and obtain a higher agreement rate with humans. However, at the same time, self-play causes model dialogues to become less diverse and does not appear to teach high-level strategic reasoning or negotiation tactics in our experiments. Although these results highlight potential room for improvement, we view them as a promising initial signal for self-play training of large language models and release all code for our environments, models, and human data collection to support future research in this area: github.com/nickatomlin/lm-selfplay.
@@ 2 Cooperative and Competitive Games
Can language model self-play be effective under both cooperative and competitive objectives? To address this question, we conducted experiments on Deal or No Deal (DoND; Lewis et al., 2017), a two-player negotiation game in which players decide how to divide a shared pool of items through natural language dialogue. Although introduced as a semi-competitive game, DoND has the special property that it can be readily adapted into either a cooperative or strictly competitive (i.e., zero-sum) game. Below, we describe the rules of DoND, how we modify its objective, and how we convert it into an environment for evaluating language models.
@@
In the original formulation, players receive a reward equal to the inner product of their value function and proposed set of objects. However, we observe that this objective can be modified to convert DoND into a cooperative game, a strictly competitive one, or anything in between. For example: if players receive rewards $R_{1}$ and $R_{2}$ in the original setting, instead setting the objective to $R_{1}+R_{2}$ for both players results in a fully cooperative game. More generally, we can define the objective for Player 1 as $R_{1}+\lambda\cdot R_{2}$ for $\lambda\in[-1,1]$, and vice versa for Player 2. In this work, we experiment with $\lambda=0$ (semi-competitive), $\lambda=1$ (cooperative), and $\lambda=-1$ (strictly competitive), although we note that in principle $\lambda$ can be tuned to smoothly interpolate between these objectives. The maximum reward that can be obtained in a single game is 10 in the strictly and semi-competitive settings and 19 in the cooperative setting. [^1] Reward is obtained by adding the item scores from each player in the cooperative setting. However, the maximum possible reward is 19 and not 20 because of the game’s constraint that at most one player can achieve the maximum item score. The maximum average self-play reward is 7.5 in the semi-competitive setting and 15 in the cooperative setting; in the zero-sum setting, average self-play scores are always zero.
@@
3 Language Model Self-Play
We begin by evaluating pretrained language models, prompted only with task instructions and the current game context, as detailed in Appendix A. Because our goal is to analyze the effects of self-play on language model behavior, we do not prompt our models with few-shot example dialogues or finetune them on task-specific data; doing so helps us avoid biasing models toward specific patterns of behavior. We then iteratively finetune these models over many rounds of self-play.
We implement a straightforward algorithm for language model self-play based on filtered behavior cloning (filtered BC; Chen et al., 2020, 2021; Zelikman et al., 2022). In this setting, two language models with identical parameters but different prompts play $K$ games and receive rewards according to their (identical) objectives. Each game produces two dialogue histories, one from each player’s perspective. The average score across games is computed and dialogues with above-average scores are kept and used for finetuning the model. This procedure is repeated for $N$ iterations or until early stopping. We set $K=500$ and $N=10$ for the majority of our experiments.
@@ 3 Language Model Self-Play
1:Input: Language model $M$, number of games per iteration $K$, number of iterations $N$, function $\operatorname{exec}$ which runs a game of self-play
@@
4:     Initialize an empty set $\mathcal{D}$
@@
12:     Compute the average $\bar{R}=\frac{1}{2K}\sum_{r\in\mathcal{R}}r$
@@
Because this approach filters out games which receive a below-average score, it requires a sufficiently capable model to “get off the ground.” In our preliminary experiments with open-weight models such as Mixtral 8x7B and LLaMA-2-70B, we found that they did not achieve enough nonzero scores to improve from the first round of self-play. In particular, weaker models struggle to follow the game instructions and routinely fail to produce compatible proposals. Therefore, we instead used GPT-3.5 (gpt-3.5-turbo-0125; Chen et al., 2021; Ouyang et al., 2022) which strikes a balance between capability and accessibility. While not an open-weight model, GPT-3.5 can still be finetuned through the OpenAI API, and it achieves sufficiently high initial scores to kickstart training. Due to the high cost of experiments, we leave investigation of other models to future work.
Figure 2: Language model self-play significantly increased model performance in both cooperative and semi-competitive games. Moreover, these results generalized to collaboration and competition with humans, leading to improvements of up to 2.5$\times$ and 6$\times$ the baseline scores, respectively. We found that human-LM baseline scores were higher in the collaborative setting as humans can help “guide” models to avoid common failure modes.
@@ 4 Human Experiments
To evaluate whether model improvements generalize beyond self-play, we built a web interface which allows humans to play DoND against our trained models. We evaluated models on both the cooperative and semi-competitive objectives across the timecourse of training. However, due to the large number of models we trained and the cost of human experiments, we only ran human evaluation on every other iteration of self-play finetuning.
@@ 5.1 Baseline Models
We conducted the majority of our experiments on GPT-3.5. Prior to self-play finetuning, GPT-3.5 performed relatively poorly, obtaining mean scores of 0.4 in the original, semi-competitive setting and 0.7 in the cooperative setting. These scores are much lower than those obtained in prior work (Lewis et al., 2017; Gandhi et al., 2023) because we did not provide models with few-shot example dialogues or task-specific finetuning data.
Qualitatively, GPT-3.5’s low scores can primarily be attributed to its inability to consistently reach complementary proposals with itself in self-play, reaching a valid agreement with its partner in only 6.8% of games across both objectives. Additionally, this baseline model relies heavily on error-handling from the environment to send messages properly and fails to remain grounded in the game’s context throughout an entire dialogue, often hallucinating new items, changes in its value function, or both.
@@
GPT-4 (Cooperative) GPT-4 (Semi-Competitive) Human (Semi-Competitive)
@@
Table 1: Mean scores of models initially finetuned on task-specific data, which was either generated using GPT-4 self-play or extracted from prior human experiments in Lewis et al. (2017). We generated finetuning data for both the cooperative and semi-competitive objectives with GPT-4, whereas pre-existing human data was only available for the semi-competitive setting. After training, models were evaluated both in self-play and via human experiments. While model scores improved in the earliest iterations of self-play finetuning, performance plateaued and even declined much earlier than in the experiments without initial training on task-specific data.
To establish a stronger baseline, we also evaluated GPT-4 using the same prompts and game environments as in the rest of our experiments, over 300 games of self-play. GPT-4 achieved a mean score of 4.3 in the semi-competitive setting and 8.8 in the cooperative setting. Qualitatively, this model was able to follow the game instructions and negotiate much more competently than GPT-3.5. Although we use GPT-4 as a reference point for model performance, we did not conduct further experiments on it due to the lack of public finetuning access at the time of experimentation.
5.2 Self-Play Finetuned Models
Improvements from self-play generalized to collaboration and competition with humans as well, with scores increasing by 6$\times$ (0.8 $\to$ 4.9) in the semi-competitive setting and 2$\times$ in the cooperative setting (4.6 $\to$ 8.7). In the cooperative setting, human-LM scores peaked at 11.6 but began to decline before the 10th iteration of self-play. We do not report scores for any model after the 10th iteration, as they tended to stabilize or even decline.
The Case of Strict Competition
For the strictly competitive setting, due to the zero-sum nature of the game, it is not informative to report mean scores in self-play (since they would always be zero). We instead evaluated the quality of trained models based on how well they performed against a separate model, GPT-4. Additionally, due to a sparsity of positive-scoring games, we modified the filtering criteria for strictly competitive self-play to also include include samples from zero-scoring games in which a valid agreement was reached. While models improved at reaching valid agreements in self-play, we found they generalized poorly against other agents. Our preliminary results indicated that even the best-performing models for this objective would routinely fail to reach agreements with humans, so we instead ran 100 games between each iteration’s model and GPT-4, confirming that the model failed to improve outside of self-play. We report results for this objective, along with further implications, extensively in Appendix D.
@@ 5.3 Comparing the Effect of Self-Play to Task-Specific Finetuning Data
We also considered the case where our initial model was finetuned on an externally provided corpus of task-specific data. For the semi-competitive objective, we finetuned models on 300 nonzero scoring games from the original human-human dataset in Lewis et al. (2017). However, because task-specific data only exists for the original task formulation, we used GPT-4 to generate a comparable amount of finetuning data for the cooperative objective; to enable fair comparison, we also used GPT-4 to generate finetuning data for the semi-competitive objective. We finetuned GPT-3.5 on the nonzero scoring games for each of these three settings and then repeated the self-play algorithm from Algorithm 1, providing a finetuned GPT-3.5 model as input instead of the baseline GPT-3.5 model.
Table 1 shows that finetuning on task-specific data, either provided by humans or distilled from a stronger model, can lead to performance improvements comparable to those obtained from language model self-play. However, the benefits of finetuning and self-play training can be combined. Running language model self-play on the finetuned models still improves model performance, but only for 1-2 iterations, after which scores begins to decline. This holds across both self-play and human experiments; moreover, we find that score improvements in self-play generally correlate with score improvements in collaboration or competition with humans.
@@ 6 Analysis
6.1 Agreements and Pareto Optimality
@@
In addition to scores increasing with each iteration of self-play finetuning, models also improved in their ability to achieve valid agreements (Table 2) and Pareto-optimal game scores (Table 3), both in self-play and in human experiments. Trained models achieved almost perfect agreement rates in self-play, with 96.4% of games ending in an agreement in the semi-competitive setting. While the agreement rates are lower in collaboration with humans, this can partially (but not wholly) be attributed to human error. For example, we observed that humans sometimes failed to read overly long messages in full, resulting in failed proposals.
@@
To understand the relationship between agreement rate and game score, we calculated Pearson correlation coefficients between completed iterations of self-play and scores achieved, before and after filtering out samples that failed to reach a valid agreement. For our self-play data under the semi-competitive objective, this yielded $\rho=0.44$ before filtering and $\rho=0.34$ after. On human-LM data with the same objective, however, the correlation drops from $\rho=0.29$ to $-0.04$ after filtering out non-scoring games. These results suggest that improvements in model performance come from an increase in agreements, which can primarily be attributed to better understanding of task instructions, following the environment rules, and not hallucinating items or proposals. Although self-play scores appear to rise after filtering out games which fail to reach a proposal, this improvement is a ruse: these improvements do not generalize to humans.
As a result, even though human-LM scores improved substantially after self-play finetuning, there is still a moderate amount of headroom on this task. More evidence for this headroom comes from Table 3, which reports the percentage of games that ended in Pareto-optimal outcomes, i.e., agreements where it would be impossible for one player to increase their score without the other player’s score decreasing. Despite improvements from self-play, models still did not obtain the maximum score in the majority of games. We expect techniques other than filtered BC may necessary to close this gap.
Semi-Competitive Cooperative
Self-Play Human Self-Play Human
Before 6.8 13.3 6.8 32.7
After 96.4 76.5 91.0 64.0
Table 2: Agreement rates (%) before and after ten rounds of self-play finetuning. Models came to more valid agreements across objectives in both self-play and human generalization. We note that the relatively lower agreement rate for human performance in the cooperative setting can be attributed to performance beginning to decline before the last iteration, as shown in Figure 2.
Semi-Competitive Cooperative
Self-Play Human Self-Play Human
Before 2.2 8.9 5.8 16.3
After 46.0 49.0 89.6 38.0
Table 3: Pareto-optimality rates (%) before and after ten rounds of self-play finetuning. Although model scores improved substantially, there is still significant headroom; models still leave many points on the table, especially in collaboration with humans.
@@
Figure 3: Mean dialogue lengths per game for every iteration of the model in semi-competitive and cooperative objectives. Dialogues under the semi-competitive objective progressively shrank in length, while dialogues under the cooperative objective grew significantly.
Figure 4: Aggregate vocabulary sizes across each model iteration’s 500 games of self-play, for both semi-competitive and cooperative objectives. In the semi-competitive setting, vocabulary size trended downward, but the model maintained and even expanded its vocabulary when trained with the cooperative objective.

### v2

@@ Abstract
Game-playing agents like AlphaGo have achieved superhuman performance through self-play, which is theoretically guaranteed to yield optimal policies in competitive games. However, most language tasks are partially or fully cooperative, so it is an open question whether techniques like self-play can effectively be used to improve language models. We empirically investigate this question in a negotiation game setting known as Deal or No Deal (DoND). Crucially, the objective in DoND can be modified to produce a fully cooperative game, a strictly competitive one, or anything in between. We finetune language models in self-play over multiple rounds of filtered behavior cloning in DoND for each of these objectives and evaluate them in self-play and in collaboration with humans. We find that language models improve substantially in self-play, achieving 14-17$\times$ higher scores in task reward after finetuning. Further, the trained models generalize to both cooperation and competition with humans, scoring 2.5-6$\times$ higher than base models. We view these results as an early promising sign for language model self-play in cooperative settings, despite a lack of theoretical guarantees.
@@ 1 Introduction
Many of the greatest achievements in artificial intelligence have occurred in two-player zero-sum (2p0s) games such as Go (Silver et al., 2016), chess (Silver et al., 2018), and heads-up poker (Brown & Sandholm, 2018). One key technique enabling these breakthroughs has been self-play, in which identical copies of a model are pitted against each other and used to generate new training data. By iteratively training on their own data from games of self-play, models like AlphaGo were able to continue improving long past the threshold of human performance. In certain types of 2p0s games, self-play is theoretically guaranteed to produce optimal policies, given sufficient model capacity and compute (Bai & Jin, 2020; Bai et al., 2020). However, in settings that involve collaboration with humans, self-play is no longer guaranteed to yield optimal policies (Strouse et al., 2021).
It is an open question whether the same principles that led to the success of models like AlphaGo can be applied to language models. Empirically, previous work on training agents to communicate via self-play has shown that they invent uninterpretable communication strategies (Kottur et al., 2017); even when initialized with natural language data, self-play can cause models to gradually diverge from human-interpretable language (Lewis et al., 2017). As a result, much work has focused on mitigating these challenges, e.g., by regularizing with models trained on human data (FAIR, 2022).
@@
We then investigate the reasons behind these improvements, finding that models trained with self-play better follow task instructions, hallucinate less, and obtain a higher agreement rate with humans. However, at the same time, self-play causes model dialogues to become less diverse and does not appear to teach high-level strategic reasoning or negotiation tactics in our experiments. Although these results highlight potential room for improvement, we view them as a promising initial signal for training language models with self-play and release all code for our environments, models, and human data collection to support future research in this area: github.com/nickatomlin/lm-selfplay.
@@ 2 Cooperative and Competitive Games
Can language model self-play be effective under both cooperative and competitive objectives? To address this question, we conducted experiments on Deal or No Deal (DoND; Lewis et al., 2017), a two-player negotiation game in which players decide how to divide a shared pool of items through natural language dialogue. Although introduced as a semi-competitive game, DoND has the special property that it can be readily adapted into either a cooperative or strictly competitive (i.e., zero-sum) game, with minimal modifications to its rules. Below, we describe the rules of DoND, how we modify its objective, and how we convert it into an environment for evaluating language models.
@@
In the original formulation, players receive a reward equal to the inner product of their value function and proposed set of objects. However, we observe that this objective can be modified to convert DoND into a cooperative game, a strictly competitive one, or anything in between. For example: if players receive rewards $R_{1}:=X$ and $R_{2}:=Y$ in the original setting, instead setting the objective to $R_{1}=R_{2}=X+Y$ for both players results in a fully cooperative game. More generally, we can define the objective for Player 1 as $R_{1}:=X+\lambda\cdot Y$ for $\lambda\in[-1,1]$, and vice versa for Player 2. In this work, we experiment with $\lambda=0$ (semi-competitive), $\lambda=1$ (cooperative), and $\lambda=-1$ (strictly competitive), although we note that in principle $\lambda$ can be tuned to smoothly interpolate between these objectives. The maximum reward that can be obtained in a single game is 10 in the strictly and semi-competitive settings and 19 in the cooperative setting. [^1] Reward is obtained by adding the item scores from each player in the cooperative setting. However, the maximum possible reward is 19 and not 20 because of the game’s constraint that at most one player can achieve the maximum item score. The maximum average self-play reward is 7.5 in the semi-competitive setting and 15 in the cooperative setting; in the zero-sum setting, average self-play scores are always zero. Finally, the average score across all Pareto-optimal outcomes in the semi-competitive setting is 6.6.
@@
1:Input: Language model $M$, number of games per iteration $K$, number of iterations $N$, function $\operatorname{exec}$ which runs a game of self-play and returns dialogues and rewards.
@@
4:     Initialize an empty set of dialogues $\mathcal{D}$
@@
12:     Compute the average reward $\bar{R}=\frac{1}{2K}\sum_{r\in\mathcal{R}}r$
@@
3 Language Model Self-Play
We begin by evaluating pretrained language models, prompted only with task instructions and the current game context, as detailed in Appendix A. In contrast to prior work, we do not prompt our models with few-shot example dialogues (Gandhi et al., 2023) or finetune them on task-specific data (Lewis et al., 2017) in the majority of our experiments. Doing so helps us avoid biasing models toward specific patterns of behavior. We then finetune these models over many rounds of self-play.
We implement a straightforward algorithm for language model self-play (Algorithm 1) based on filtered behavior cloning (filtered BC; Chen et al., 2020; 2021; Zelikman et al., 2022). In this setting, two language models with identical parameters but different prompts play $K$ games and receive rewards according to their (identical) objectives. Each game produces two dialogue histories, one from each player’s perspective. The average score across games is computed, and dialogues with above-average scores are kept and used for finetuning the model. This procedure is repeated for $N$ iterations or until early stopping. We set $K=500$ and $N=10$ for the majority of our experiments.
We ran experiments on GPT-3.5 (gpt-3.5-turbo-0125; Ouyang et al., 2022), using the OpenAI API for finetuning. At the time of experimentation, GPT-3.5 was the most capable model for which finetuning was publicly available. We also ran preliminary experiments with open-weight models such as Mixtral 8x7B and LLaMA-2-70B, but we found that they did not achieve enough nonzero scores to improve from the first round of self-play. Finally, we ran a small-scale baseline experiment on GPT-4 (gpt-4-turbo-2024-04-09; OpenAI, 2023). Due to the high cost of experiments, we leave investigation of other models to future work.
@@ 4 Human Experiments
To evaluate whether model improvements generalize beyond self-play, we built a web interface which allows humans to play DoND against our trained models. We evaluated models on both the cooperative and semi-competitive objectives across the timecourse of training. As discussed in Section 5.2, we did not run human experiments on the strictly competitive objective, where models perform poorly. Additionally, due to the large number of models we trained and the cost of human experiments, we only ran human evaluation on every other iteration of self-play finetuning.
@@
Figure 2: Language model self-play significantly increased model performance in both cooperative and semi-competitive games. Moreover, these results generalized to collaboration and competition with humans, leading to improvements of up to 2.5$\times$ and 6$\times$ the baseline scores, respectively. We found that human-LM baseline scores were higher in the cooperative setting as humans can help “guide” models to avoid common failure modes.
@@ 5.1 Baseline Models
We first evaluated language models without any self-play training. Because we did not provide few-shot example dialogues, GPT-3.5 performed relatively poorly, obtaining a mean score of 0.4 in the original, semi-competitive setting and 0.7 in the cooperative setting. In contrast, GPT-4 achieved much higher mean scores of 4.3 in the semi-competitive setting and 8.8 in the cooperative setting. Although we use GPT-4 as a reference point for model performance, we did not conduct further experiments on it due to the lack of public finetuning access at the time of experimentation.
GPT-3.5’s low scores can primarily be attributed to its inability to consistently reach complementary proposals with itself in self-play, reaching a valid agreement in only 6.8% of games across both objectives. Additionally, this baseline model relies heavily on error-handling from the environment to send messages properly and fails to remain grounded in the game’s context throughout an entire dialogue, often hallucinating new items, changes in its value function, or both.
@@
5.2 Self-Play Finetuned Models
Improvements from self-play generalized to collaboration and competition with humans as well, with scores increasing by 6$\times$ in the semi-competitive setting (0.8 $\to$ 4.9) and 2.5$\times$ in the cooperative setting (4.6 $\to$ 11.6). In the cooperative setting, we note that human-LM scores peaked at 11.6, but began to decline before the 10th iteration of self-play. We do not report scores for any model after the 10th iteration, as they tended to stabilize or even decline.
The Case of Strict Competition
Between fully rational agents, communication is not useful in a 2p0s game (Crawford & Sobel, 1982). Additionally, for the strictly competitive setting, due to the zero-sum nature of the game, it is not informative to report mean scores in self-play, as the average score in this setting will always be zero. We instead evaluated the quality of trained models based on how well they performed against a separate model, GPT-4. Additionally, due to a sparsity of positive-scoring games, we modified the filtering criteria for strictly competitive self-play to also include samples from zero-scoring games in which a valid agreement was reached. While models improved at reaching valid agreements in self-play, we found they generalized poorly against other agents. Our preliminary results indicated that even the best-performing models for this objective would routinely fail to reach agreements with humans, so we instead ran 100 games between each iteration’s model and GPT-4, confirming that the model failed to improve outside of self-play.
We report results for this objective, along with further implications, extensively in Appendix D.
GPT-4 (Cooperative) GPT-4 (Semi-Comp.) Human (Semi-Comp.)
@@ 5.2 Self-Play Finetuned Models
Table 1: Mean scores of models initially finetuned on task-specific data, which was either generated using GPT-4 self-play or extracted from prior human experiments in Lewis et al. (2017). After training, models were evaluated both in self-play and via human experiments. While model scores improved in the earliest iterations of self-play finetuning, performance plateaued and even declined much earlier than in the experiments without initial training on task-specific data.
@@ 5.3 Comparing the Effect of Self-Play to Task-Specific Finetuning Data
We also considered the case where our initial model was finetuned on an externally provided corpus of task-specific data. For the semi-competitive objective, we finetuned models on 300 nonzero scoring games from the original human-human dataset in Lewis et al. (2017). However, because task-specific data only exists for the original task formulation, we used GPT-4 to generate a comparable amount of finetuning data for the cooperative objective; to enable fair comparison, we also used GPT-4 to generate finetuning data for the semi-competitive objective. We finetuned GPT-3.5 on the nonzero scoring games for each of these three settings and then repeated the self-play algorithm from Algorithm 1, providing a finetuned GPT-3.5 model as input instead of the baseline GPT-3.5.
We found that finetuning drastically improves the performance of base models, as shown in Table 1. Applying self-play training on top of finetuned models provided slight additional gains, ranging from +6% (11.0 $\to$ 11.7) in the cooperative setting to +38% (4.2 $\to$ 5.8) in the semi-competitive setting. These improvements are much smaller than the improvements from self-play finetuning on base models. However, we note that: (a) our finetuning process may be viewed as a single round of filtered BC where zero-scoring games are filtered out, (b) self-play provides further gains on top of task-specific finetuning, and (c) task-specific finetuning data may not be available for all tasks. Therefore, we view finetuning not as a replacement for self-play, but as a complementary method.
Agreement Pareto-Optimality
Semi-Competitive Cooperative Semi-Competitive Cooperative
Self-Play Human Self-Play Human Self-Play Human Self-Play Human
Before 6.8 13.3 6.8 32.7 2.2 8.9 5.8 16.3
After 96.4 76.5 91.0 64.0 46.0 49.0 89.6 38.0
Table 2: Agreement and Pareto-optimality rates (%) before and after ten rounds of self-play finetuning. Models show significant improvements across objectives in both self-play and human generalization. We observe a relatively lower agreement rate for human performance in the cooperative setting, as well as the remaining headroom in Pareto-optimality, especially in human collaboration.
@@ 6 Analysis
6.1 Errors, Agreements, and Pareto Optimality
In order to understand the effectiveness of language model self-play, we analyzed the frequency of errors and successful agreements over the course of training. We also analyzed the rate of Pareto-optimal outcomes, i.e., those where neither player’s score can improve without reducing the other’s.
Prior to self-play finetuning, GPT-3.5 frequently made errors such as submitting invalid proposals or sending messages and proposals at the same time. When this happened, the game environment would provide an error message, as detailed in Appendix A, and the model would be given another chance to generate a message. With the baseline model, errors were generated in 14% of semi-competitive games and 56% of cooperative games. If the model received five error messages in a row, the game aborted, and both players received a score of zero; this outcome occurred in just 1% of semi-competitive games but in 22% of cooperative games. After self-play finetuning, errors occurred in less than 1% of games, suggesting that models effectively learned the game rules. We present additional results on error and abort frequencies in Appendix E.
@@ 6.1 Errors, Agreements, and Pareto Optimality
Models also improved in their ability to achieve valid agreements and Pareto-optimal game scores, both in self-play and in human experiments, as shown in Table 2. Trained models achieved almost perfect agreement rates in self-play, with 96.4% of games ending in an agreement in the semi-competitive setting. While the agreement rates are lower in collaboration with humans, this can partially (but not wholly) be attributed to human error. For example, we observed that humans sometimes failed to read overly long messages in full, resulting in failed proposals; however, such errors may also be viewed as the fault of models failing to communicate efficiently.
@@
We found that model improvements after self-play can primarily be attributed to an increased rate of agreement. To justify this claim, we calculated Pearson correlation coefficients between completed iterations of self-play and scores achieved, before and after filtering out samples that failed to reach a valid agreement. For our self-play data under the semi-competitive objective, this yielded $\rho=0.44$ before filtering and $\rho=0.34$ after. On human-LM data with the same objective, however, the correlation drops from $\rho=0.29$ to $-0.04$ after filtering out non-scoring games. In other words, although self-play scores appear to rise after filtering out games which fail to reach a proposal, these improvements do not generalize to humans.
We hypothesize that the increased agreement rate can primarily be attributed to better understanding of task instructions, following the environment rules, and not hallucinating items or proposals, rather than the acquisition of strategic negotiation or optimization behavior. As further evidence for this claim, we note that even after self-play finetuning, models routinely missed opportunities for better scores: only 49% of semi-competitive games with humans resulted in Pareto-optimal outcomes, and only 38% of cooperative games did so. As a result, we note that there is still substantial headroom on this task, although we expect techniques other than filtered BC may necessary to close this gap.
@@
Figure 3: Mean dialogue lengths (left) and aggregate vocabulary sizes (right) for every model iteration, for both semi-competitive and cooperative objectives. Dialogues under the semi-competitive objective progressively shrank in length, while dialogues under the cooperative objective grew significantly longer. Similarly, in the semi-competitive setting, vocabulary size trended downward, but the model maintained and even expanded its vocabulary when trained with the cooperative objective.